# Decoding Visual Experience and Mapping Semantics through Whole-Brain Analysis Using fMRI Foundation Models

### v2

@@ 2.2 Post-hoc analysis of Whole Brain Visual Experience
To showcase our model’s capability in decoding visual experiences, we performed an extensive post-hoc analysis of whole-brain visual experiences using our WAVE model. Employing a self-supervised approach, we clustered original images into five distinct groups and analyzed the corresponding patterns of brain activity. As illustrated in Fig. 3, the brain saliency maps highlight the regions of interest (ROIs) activated by each image cluster. These findings reveal that different clusters are associated with distinct neural activation patterns, but also demonstrate that similar brain networks were important for decoding across different clusters.
Figure 3: Visualization of Post-Hoc Whole-Brain Visual Clustering Analysis. a, t-SNE visualization of image embeddings, differentiated by five clusters identified via K-Means, each represented in unique colors. b-f, Each subfigure presents word clouds generated from image labels, with word sizes proportional to their frequency of occurrence within the cluster. The subfigure titles, generated by entering the words into ChatGPT-4, summarize the thematic essence of each cluster. Accompanying each word cloud are selected image samples and a whole-brain saliency map highlighting the top 20 regions of interest relevant to the cluster. g, Quantitative analysis showing the distribution of regions across the Yeo 7 networks for each of the five clusters, providing insights into the network-based localization of visual processing associated with different categories.
@@ 2.3 Zero-Shot Human Imagination Decoding
Figure 4: WAVE Zero-Shot brain imagination decoding results. a, Scenario-based scale analysis utilizing two-way-identification accuracy. Using a separate dataset (N=24 younger adults imagining different scenarios during fMRI), this analysis measures the cosine distance of each scenario’s features to those in the training dataset (BOLD5000). The scatter plot shows the correlation of these cosine distances with decoding accuracy: scenarios that are more similar to the training dataset had more accurate predictions. b, Example of WAVE model zero-shot reconstructed images from imagination recording fMRI sessions. The text stimuli describing the museum scenario are presented above the example image, illustrating the model’s capability to generate visual reconstructions based on described scenarios.
@@ 3 Discussion
In this paper, we presented a novel approach to decode visual representations present throughout the brain by leveraging a pre-trained foundation model. Compared to the state-of-the-art approaches to decoding visual processing, we found that our model improved on predictive accuracy particularly when assessing the higher-level visual and semantic similarity between reconstructed and visual images. Saliency maps visualized regions most important to these predictions and identified predominantly regions in higher-order brain networks such as the frontoparietal and default mode networks, suggesting that higher-order processing was critical to this decoding. In line with this, masking the visual network still produced high levels of accuracy on the same visual decoding task, although it did lead to reductions in accuracy as would be expected. Ablating each network in turn showed that after the visual network, the dorsal attention and default mode networks contributed the most to predictions. This is in line with previous research: the dorsal attention network is adjacent to the visual network and performs higher level visual functions. The default mode network has a key role in semantic processing and “making sense” [64] of visual input, as shown by the presence of a retinotopic code in this network [51]. Post-hoc analyses comparing our model to others suggest that our model was better able to model the semantic information in images due to improved identification of key semantic features, sometimes with an associated loss of accuracy in lower-level visual similarity such as pixel correlation. A scaling analysis in a separate dataset that involved participants making semantic verbal descriptions of scenarios that they imagined during fMRI showed that decoding was improved for scenarios that were more semantically similar to images in the training dataset (e.g., “internet” or “movie” scenarios vs. worse performance for “barbecue” or “resting” scenarios). Overall, these findings highlight that foundation models of fMRI data allow for improved decoding of complex higher-level visual and semantic features in images even when applied to relatively small, specialized datasets that have predominantly traditionally been used to understand lower-level visual processing. We also show that essential information for visual decoding is present in non-visual, higher-order brain networks, particuarly the dorsal attention and default mode networks, further emphasizing the role of non-visual regions in visual processing. Our scaling analysis further suggests the importance of semantic information in the performance of this model.
@@
In addition to our primary finding of improved performance over other state-of-the-art approaches, there are several important implications of our work for neuroscience. The ability to improve performance on visual processing using a foundation model trained and applied to the whole brain adds to recent literature highlighting the importance of non-visual regions to visual processing. For example, a recent study identified a retinotopic code in the default mode network that is hypothesized to provide a mapping between past experiences and incoming visual information [51]. This is in line with ideas from predictive coding [1] and active inference [15] theories of the brain that propose incoming visual information is continuously compared to an internal model of the world provided via top-down connections between visual and higher-order brain regions. The default mode network has been proposed to be critical in this “sense-making” [64], in line with our finding that it contributed more than other non-visual networks to improved predictions. Related work suggests that visual images may be encoded in non-visual regions in terms of the actions they afford  [56], in line with theories that consider the fundamental role of the brain to induce actions and predict their consequences as a means of increasing the survival of the individual [7]. Conceptually, our work builds on this by highlighting that higher-order brain networks, particuarly the default mode network, contribute critical information to visual decoding and allow for improved separation of visually similar images into distinct clusters based on higher-order visual and semantic information: for example, the separation of wild dogs from domestic dogs.
@@
Our post-hoc explainability analysis also provides some information about exactly which brain regions are contributing to decoding different categories of images. While it is difficult to interpret exactly what certain brain activity patterns mean for the information they represent without engaging in overly speculative reverse inference [40], these data do add important interpretability to our findings in several ways. Breaking down these activation patterns by large-scale brain network, regions in the default mode network were important for all five clusters, in line with the role of this network in semantic processing [6] and matching the results of our network-ablation anaylsis. The frontoparietal and dorsal attention networks were also important for predictions in all five clusters: these networks are involved in domain-general task-related processing and higher-order visual processing [33, 60] and this analysis suggests they contain information for decoding visual information in various domains. Some clusters did show unique ROIs: limbic regions were only involved in predictions of pets, potentially reflecting the particularly emotional nature of information processing of these images. Sensorimotor regions were most important for decoding the cluster representing wild animals, as shown in Fig. 3g, suggesting a possible additional engagement of areas involved in spatial awareness and movement perception [19] when processing these images. This analysis further emphasizes the importance of whole-brain non-visual regions for visual processing and further suggests that certain images may engage different distributed neural processes depending on their contents.
@@
The final important implication of this work is methodological. We believe that our approach allows for significant further advances in understanding the role of non-visual regions in visual processing by facilitating the mapping of these higher-order processes using foundation models. Moreover, we believe that these models hold significant potential in advancing our understanding of brain-behavior relationships particularly for more complex traits that vary more across people. These models allow us to take advantage of the large amounts of non-specific fMRI data that now exist and may help us improve our predictions of these more complex brain-behavior relationships, beyond functional decoding and into more trait-level brain-behavior analyses. This has huge implications for the field, as these tasks have previously required the collection of either highly specific data or more general data in thousands of individuals [34]. This has led to significant consternation in the field regarding the practicality of using MRI to understand behavior and led some researchers to suggest that data for understanding brain-behavior relationships for complex traits should only be done by consortia [18]. Using similar principles to those proposed by other researchers [22], we believe that foundation models may boost the predictive power of models that aim to predict more variable brain-behavior relationships, reducing the sample size required for individual labs and providing a new ”middle path” to robust brain-behavior predictions.
@@ 4.1 Dataset
For our study, we utilized the BOLD5000 dataset, a publicly available fMRI image collection [9]. This dataset comprises fMRI recordings captured as subjects viewed a series of images. The BOLD5000 dataset contains 4,916 unique images, of which 2,000 are sourced from the COCO dataset [32], and 1,916 images are from ImageNet [12]. In each trial, subjects were exposed to each image as a visual stimulus. Among these, 4,803 images were presented once per trial, while 113 images were shown multiple times, either three or four times across several trials, resulting in a total of 5,254 stimulus trials. The dataset documents responses from four subjects, three of whom completed all 5,254 recording trials. In our analysis, we focus on the data from all four subjects who completed the full or partial set of trials.
We adopt the standardized train/test set split employed in existing BOLD5000 image reconstruction studies, as outlined in Chen et al. [10]. For the training set of each subject, we utilized trials with non-repeated image stimuli, encompassing a total of 4,803 samples. For the test set, we consolidated repeated image stimulus trials, resulting in 113 test samples. Furthermore, BOLD5000 provides labels for each image, which we incorporated as the text modality in our analysis.
@@
Additionally, we employed a validation dataset from fMRI imagination recordings as per [3]. This dataset includes 20 different scenarios such as resting, dancing, and museum visits. Subjects provided varied text descriptions for these scenarios. The fMRI data were recorded as subjects imagined themselves in the scenarios prompted by the text descriptions they had written. Each scenario was repeated five times during the fMRI sessions. In our zero-shot analysis, we averaged the fMRI signals from the repeated scenarios to create a single representative signal for each scenario. We engaged 24 young subjects for this study, resulting in a total of $24\times 20=480$ imagination fMRI clips accompanied by text stimuli.
@@
In this paper, we utilize a foundational fMRI model incorporating several large-scale fMRI datasets for pre-training [55]. These include the Human Connectome Project [57] and multiple datasets available on OpenNeuro [35], comprising a total of 11,980 fMRI runs from 1,726 individuals across 34 datasets. Additionally, we employed a pre-trained image classifier to evaluate semantic accuracy metrics. This classifier was originally trained on ImageNet1K [47], a dataset that includes approximately 1.2 million images distributed over 1,000 categories, each representing a distinct object or concept.
@@
After the preprocessing steps, we parcellated each preprocessed fMRI run using the DiFuMo atlas [11], normalizing the individual network time courses to have a mean of zero and unit variance. We opted for the DiFuMo atlas with 1024 dimensions, effectively transforming each volume $V_{t}\in\mathbb{R}^{x\times y\times z}$ into a vectorized form $X_{t}\in\mathbb{R}^{1024}$ as shown in Fig. 5.
In the subsequent step of segmenting BOLD clips, we took into account the hemodynamic effects on BOLD signals[39]. To accommodate the delay introduced by these hemodynamic effects, we adjusted our segmentation to account for a one TR delay. In our dataset configuration, each segmented BOLD clip is characterized by five TRs, resulting in the final fMRI input for our model being $X\in\mathbb{R}^{5\times 1024}$, where $X$ represents our input data.
@@ 4.2 WAVE Model
Our model employs state-of-the-art deep learning [31, 17] and foundation model [5] techniques, structured into two main phases: contrastive learning [43] and diffusion model training [45]. Unlike approaches that merge semi-supervised training methods, our strategy benefits from progressive learning [48]. We initiate with contrastive learning to refine initial insights, followed by the diffusion model phase for enhanced training. Decoding fMRI data poses significant challenges due to its complexity. To address this, we incorporate a pre-trained large fMRI foundation model[55]. In the contrastive learning stage, we integrate image labels as an additional modality, recognizing that label data alone may not provide comprehensive information. To bridge this gap, we employ prompt learning to enrich the text encoder with supplementary image and fMRI details. For the diffusion model training, we introduce a novel approach by employing a diffusion prior[44, 49] to transform fMRI latent representations into image latent representations. This process is completed with a Pre-trained Versatile Diffusion Image Variation Decoder[61], which translates these image latent representations back into visual images. This methodology not only leverages advanced machine learning techniques but also introduces an innovative integration of fMRI data into the image generation process, enhancing the model’s ability to interpret and visualize complex neuroimaging datasets.
@@
In our research, we adopted a pre-trained Foundation model [55] to serve as the fMRI encoder. This model, structured based on the GPT-2 framework [42], processes the preprocessed fMRI data into fMRI latent representations. As specified in Section 4.1, the input data $X$, is conformed to the preprocessed structure $X\in\mathbb{R}^{t\times 1024}$, where 1024 denotes the parcellated brain networks. The fMRI Foundation model employs self-supervised learning (SSL) on extensive fMRI datasets, including the Human Connectome Project [57] and datasets available through OpenNeuro [35]. It assimilates both the sequential temporal fMRI information and TR embeddings, which function similarly to position embeddings [58] to indicate the temporal intervals between fMRI sequences. The self-supervised learning task is to predict the subsequent timepoint sequence given an input sequence $X_{t}$; thus, the model predicts $X_{t+1}$.
In this study, we processed an fMRI run, retaining the spatial and temporal features to form the input $X$, where $X\in\mathbb{R}^{5\times 1024}$. Each fMRI segment has 5 TRs. The attention is on the embedded fMRI sequence. Based on the SSL paradigm, we took the predicted masked token, which is the output $O_{t}\in\mathbb{R}^{768}$ where $t=6$. The output $O_{6}$ was taken as our encoded fMRI latent features.
@@
In this study, we employ multi-modal contrastive learning, capitalizing on the integration of diverse data forms to enrich the feature representation of fMRI scans in a latent space. Our approach is distinct in that it encompasses not only visual data but also textual information through the inclusion of image labels. This incorporation aims to augment the semantic understanding within our model. To facilitate this, we utilize a pre-trained Vision Transformer (ViT) [13] based CLIP model [43], maintaining the integrity of its parameters by freezing them throughout the training phase.
@@
$\mathcal{L_{\text{CLIP}}}(z_{1},z_{2})=-\frac{1}{2|B|}\sum_{i=1}^{|B|}\left[\log\frac{\exp({z}_{\text{1}_{i}}\cdot{z}_{\text{2}_{i}}/\tau)}{\sum_{j=1}^{|B|}\exp({z}_{\text{1}_{i}}\cdot{z}_{\text{2}_{j}}/\tau)}+\log\frac{\exp({z}_{\text{2}_{i}}\cdot{z}_{\text{1}_{i}}/\tau)}{\sum_{j=1}^{|B|}\exp({z}_{\text{2}_{i}}\cdot{z}_{\text{1}_{j}}/\tau)}\right],$ (1)
@@
Furthermore, we assess the efficacy of our training process through contrastive learning accuracy metrics. Detailed explanations of these methodologies and the results are provided in the Appendix.
@@
To enhance our contrastive learning framework, we were inspired by prompt learning [65, 27, 62] in the AI field, which is particularly effective when text labels serve as one modality in contrastive learning. Text labels, while compact, inherently contain less information than images. Prompt learning, a technique tailored for text encoders, enriches this modality by generating context that extracts more nuanced information from the text. Considering the fMRI, image, and text share the same information. We introduced a lightweight architecture called Meta-Net designed to extract features from fMRI $\mathcal{H_{\textit{fMRI}}}$ and image $\mathcal{H_{\textit{image}}}$ latent space, $\mathcal{H}\in\mathbb{R}^{768}$, and $\text{features}~{}\mathcal{F}\in R^{1536}$ and output prompt learnable vectors $V\in\mathbb{R}^{8}$ for the text encoder. The vectors will be concatenated with tokenized text together input to the CLIP text encoder. Unlike CLIP text and image encoders, the gradients of the Meta-Net were not frozen during the contrastive training, where the Meta-Net was continuously trained and learnable.

### v3

@@ 2.2 Post-hoc analysis of Whole Brain Visual Experience
To investigate how distinct visual concepts recruit specific whole-brain networks, we performed an extensive post-hoc semantic profiling analysis of whole-brain visual experiences using our WAVE model. We applied unsupervised clustering to the stimulus space, partitioning original images into five distinct semantic groups and analyzed the corresponding patterns of brain activity. As illustrated in Fig. 3, the brain saliency maps highlight the regions of interest (ROIs) activated by each image cluster. These findings reveal that different clusters are associated with distinct neural activation patterns, but also demonstrate that similar brain networks were important for decoding across different clusters.
Figure 3: Semantic Profiling of Whole-Brain Visual Representations. a, t-SNE projection of image embeddings, colored by five distinct semantic clusters identified via K-Means. b-f, Detailed profiles for each cluster. Word clouds illustrate the most frequent object labels within each group The subfigure titles, generated by entering the words into ChatGPT-4, summarize the thematic essence of each cluster. Accompanying each word cloud are selected image samples and a whole-brain saliency map highlighting the top 20 regions of interest relevant to the cluster. g, Network-level decomposition of the saliency maps. The bar chart quantifies quantifies the distribution of top regions across Yeo-7, providing insights into the network-based localization of visual processing associated with different categories.
@@ 2.3 Zero-Shot Human Imagination Decoding
Figure 4: Generalization to Zero-Shot Mental Imagery. a, WAVE performance on an independent dataset ($N=24$), where participants imagined scenarios based on verbal prompts. This analysis measures the cosine distance of each scenario’s features to those in the training dataset (BOLD5000). The scatter plot shows the correlation of these cosine distances with decoding accuracy: scenarios that are more similar to the training dataset had more accurate predictions. b, Example of WAVE model zero-shot reconstructed images from imagination recording fMRI sessions. The text stimuli describing the museum scenario are presented above the example image, illustrating the model’s capability to generate visual reconstructions based on described scenarios.
@@ 3 Discussion
In this paper, we presented a novel approach to decode visual representations present throughout the brain by leveraging a pre-trained foundation model. Compared to the state-of-the-art approaches to decoding visual processing, we found that our model improved on predictive accuracy particularly when assessing the higher-level visual and semantic similarity between reconstructed and visual images. Saliency maps visualized regions most important to these predictions and identified predominantly regions in higher-order brain networks such as the frontoparietal and default mode networks, suggesting that higher-order processing was critical to this decoding. In line with this, masking the visual network still produced high (although reduced) levels of accuracy on the same visual decoding task. Ablating each network in turn showed that after the visual network, the dorsal attention and default mode networks contributed the most to predictions. This is in line with previous research: the dorsal attention network is adjacent to the visual network and performs higher level visual functions. The default mode network has a key role in semantic processing and “making sense” Yeshurun et al. [2021] of visual input, as shown by the presence of a retinotopic code in this network [Steel et al., 2024]. Post-hoc analyses comparing our model to others suggest that our model was better able to model the semantic information in images due to improved identification of key semantic features, sometimes with an associated loss of accuracy in lower-level visual similarity such as pixel correlation. Furthermore, our scaling analysis on an independent imagination dataset demonstrates the model’s capacity for zero-shot generalization. We observed that decoding accuracy was modulated by semantic alignment: scenarios conceptually closer to the training distribution (e.g., “internet” or “movie”) were reconstructed with higher fidelity than distinct concepts (e.g., “barbecue”). This gradient in performance highlights that the model relies on shared semantic representations rather than memorized visual patterns. Crucially, these results validate the utility of foundation models for whole-brain analysis. Whole-brain fMRI is characterized by a massive feature space, high noise-to-signal ratio, and complex non-linear challenges where standard linear models or MLPs often struggle to converge. By leveraging pre-trained temporal and spatial priors, our approach effectively extracts signal from this high-dimensional noise, unlocking the rich semantic information encoded in non-visual networks like the dorsal attention and default mode networks, even when applied to datasets with very few individuals.
@@
In addition to our primary finding of improved performance over other state-of-the-art approaches, there are several important implications of our work for neuroscience. The ability to improve performance on visual processing using a foundation model trained and applied to the whole brain adds to recent literature highlighting the importance of non-visual regions to visual processing. For example, a recent study identified a retinotopic code in the default mode network that is hypothesized to provide a mapping between past experiences and incoming visual information [Steel et al., 2024]. This is in line with ideas from predictive coding [Aitchison and Lengyel, 2017] and active inference [Friston et al., 2009] theories of the brain that propose incoming visual information is continuously compared to an internal model of the world provided via top-down connections between visual and higher-order brain regions. The default mode network has been proposed to be critical in this “sense-making” Yeshurun et al. [2021], in line with our finding that it contributed more than other non-visual networks to improved predictions. Related work suggests that visual images may be encoded in non-visual regions in terms of the actions they afford  [Thornton and Tamir, 2024], in line with theories that consider the fundamental role of the brain to induce actions and predict their consequences as a means of increasing the survival of the individual [Buzsaki, 2019]. Conceptually, our work builds on this by highlighting that higher-order brain networks, particuarly the default mode network, contribute critical information to visual decoding and allow for improved separation of visually similar images into distinct clusters based on higher-order visual and semantic information: for example, the separation of wild dogs from domestic dogs.
@@
Our post-hoc explainability analysis also provides some information about exactly which brain regions are contributing to decoding different categories of images. While it is difficult to interpret exactly what certain brain activity patterns mean for the information they represent without engaging in overly speculative reverse inference [Poldrack, 2011], these data do add important interpretability to our findings in several ways. Breaking down these activation patterns by large-scale brain network, regions in the default mode network were important for all five clusters, in line with the role of this network in semantic processing [Brandman et al., 2021] and matching the results of our network-ablation anaylsis. The frontoparietal and dorsal attention networks were also important for predictions in all five clusters: these networks are involved in domain-general task-related processing and higher-order visual processing [Marek and Dosenbach, 2018, Vossel et al., 2014] and this analysis suggests they contain information for decoding visual information in various domains. Some clusters did show unique ROIs: limbic regions were only involved in predictions of pets, potentially reflecting the particularly emotional nature of information processing of these images. Sensorimotor regions were most important for decoding the cluster representing wild animals, as shown in Fig. 3g, suggesting a possible additional engagement of areas involved in spatial awareness and movement perception [Grosbras et al., 2012] when processing these images. This analysis further emphasizes the importance of whole-brain non-visual regions for visual processing and further suggests that certain images may engage different distributed neural processes depending on their contents.
We acknowledge several limitations in the current study. First, while our generative reconstructions achieve high semantic accuracy, we caution that diffusion models can introduce “hallucinated” high-frequency details based on their training priors. Therefore, the pixel-wise accuracy of our reconstructions should be interpreted as a synthesis of neural decoding and generative interpolation. Second, our claims regarding the ”causal” contribution of the DMN are based on computational lesioning (masking); future work using transcranial magnetic stimulation (TMS) or patient studies would be required to establish true biological causality.
@@
The final important implication of this work is methodological. We believe that our approach allows for significant further advances in understanding the role of non-visual regions in visual processing by transferring learned priors from vast, non-specific fMRI repositories to specialized tasks. Moreover, this capability holds particular promise for advancing brain-behavior analyses, especially for complex traits that exhibit high inter-subject variability. These models allow us to take advantage of the large amounts of non-specific fMRI data that now exist and may help us improve our predictions of these more complex brain-behavior relationships, beyond functional decoding and into more trait-level brain-behavior analyses. This has huge implications for the field, as these tasks have previously required the collection of either highly specific data or more general data in thousands of individuals [Marek et al., 2022]. This has led to significant consternation in the field regarding the practicality of using fMRI to understand behavior and led some researchers to suggest that data for understanding brain-behavior relationships for complex traits should only be done by consortia [Gratton et al., 2022]. Using similar principles to those proposed by other researchers [He et al., 2022], brain foundation model captured universal neural activities can significantly boost predictive power in smaller datasets. Consequently, this approach has the potential to democratize robust brain-behavior research, allowing for precise modeling of complex cognitive traits without the prohibitive requirement of massive sample sizes.
@@ 4.1 Dataset
For our study, we utilized the BOLD5000 dataset, a publicly available fMRI image collection [Chang et al., 2019]. This dataset comprises fMRI recordings captured as subjects viewed a series of images. The BOLD5000 dataset contains 4,916 unique images, of which 2,000 are sourced from the COCO dataset [Lin et al., 2014], and 1,916 images are from ImageNet [Deng et al., 2009]. In each trial, subjects were exposed to each image as a visual stimulus. Among these, 4,803 images were presented once per trial, while 113 images were shown multiple times, either three or four times across several trials, resulting in a total of 5,254 stimulus trials. The dataset documents responses from four subjects, three of whom completed all 5,254 recording trials. In our analysis, we focus on the data from all four subjects who completed the full or partial set of trials.
We selected BOLD5000 to capture the comprehensive dynamics of high-level visual experience. While widely used repositories like the Natural Scenes Dataset (NSD) Allen et al. [2022] prioritize high-throughput sampling via rapid presentation (3s stimulus, 1s gap), such rapid pacing effectively truncates the temporal window necessary for slow, top-down cognitive processes to fully unfold. On the other hand, BOLD5000 employs a slow event-related design with extended inter-stimulus intervals (9s gap). This extended duration allows for the complete emergence of delayed feedback signals and preserves the rich temporal dynamics required by our pretrained autoregressive fMRI encoder Thomas et al. [2022]. This enables us to decode higher-order semantics from whole-brain activity, capturing the full scope of the visual experience.
We adopt the standard train/test split employed in existing BOLD5000 image reconstruction studies, as outlined in Chen et al. [2023]. For the training set of each subject, we utilized trials with non-repeated image stimuli, encompassing a total of 4,803 samples. For the test set, we consolidated repeated image stimulus trials, resulting in 113 test samples. Furthermore, BOLD5000 provides labels for each image, which we incorporated as the text modality in our analysis.
@@
Additionally, we employed a validation dataset from fMRI imagination recordings as per Anderson et al. [2020]. This dataset includes 20 different scenarios such as resting, dancing, and museum visits. Subjects provided varied text descriptions for these scenarios. The fMRI data were recorded as subjects imagined themselves in the scenarios prompted by the text descriptions they had written. Each scenario was repeated five times during the fMRI sessions. In our zero-shot analysis, we averaged the fMRI signals from the repeated scenarios to create a single representative signal for each scenario. We engaged 24 young subjects for this study, resulting in a total of $24\times 20=480$ imagination fMRI clips accompanied by text stimuli.
@@
In this paper, we utilize a foundational fMRI model incorporating several large-scale fMRI datasets for pre-training [Thomas et al., 2022]. These include the Human Connectome Project [Van Essen et al., 2013] and multiple datasets available on OpenNeuro [Markiewicz et al., 2021], comprising a total of 11,980 fMRI runs from 1,726 individuals across 34 datasets. Crucially, the pre-training curriculum includes not only resting-state acquisitions but also a wide array of active tasks, such as passive visual perception (e.g., movie-watching), working memory recall, language processing, and motor coordination. This exposure to varied tasks allows the model to learn generalized spatiotemporal representations that transcend specific experimental paradigms. Additionally, we employed a pre-trained image classifier to evaluate semantic accuracy metrics. This classifier was originally trained on ImageNet1K [Russakovsky et al., 2015], a dataset that includes approximately 1.2 million images distributed over 1,000 categories, each representing a distinct object or concept.
@@
After the preprocessing steps, we parcellated each preprocessed fMRI run using the DiFuMo atlas [Dadi et al., 2020], normalizing the individual network time courses to have a mean of zero and unit variance. We opted for the DiFuMo atlas with 1024 dimensions, effectively transforming each volume $V_{t}\in\mathbb{R}^{x\times y\times z}$ into a vectorized form $X_{t}\in\mathbb{R}^{1024}$ as shown in Fig. 5.
In the subsequent step of segmenting BOLD clips, we took into account the hemodynamic effects on BOLD signals[Pinsky and Payen, 2005]. To accommodate the delay introduced by these hemodynamic effects, we adjusted our segmentation to account for a one TR delay. In our dataset configuration, each segmented BOLD clip is characterized by five TRs, resulting in the final fMRI input for our model being $X\in\mathbb{R}^{5\times 1024}$, where $X$ represents our input data.
@@ 4.2 WAVE Model
Our model employs state-of-the-art deep learning [LeCun et al., 2015, Goodfellow et al., 2016] and foundation model Bommasani et al. [2021] techniques, structured into two main phases: contrastive learning [Radford et al., 2021] and diffusion model training [Rombach et al., 2022]. Unlike approaches that rely on hybrid training schemes Scotti et al. [2024], our framework leverages a progressive learning framework for better fMRI foundation model adaption [Rusu et al., 2016]. We initiate with contrastive learning to refine initial insights, followed by the diffusion model phase for enhanced training. Decoding fMRI data poses significant challenges due to its complexity. To address this, we incorporate a pre-trained large fMRI foundation model[Thomas et al., 2022]. In the contrastive learning stage, we integrate image labels as an additional modality, recognizing that label data alone may not provide comprehensive information. To bridge this gap, we employ prompt learning to enrich the text encoder with supplementary image and fMRI details. For the diffusion model training, we introduce a novel approach by employing a diffusion prior[Ramesh et al., 2022, Scotti et al., 2024] to transform fMRI latent representations into image latent representations. This process is completed with a Pre-trained Versatile Diffusion Image Variation Decoder[Xu et al., 2023], which translates these image latent representations back into visual images. This methodology not only leverages advanced machine learning techniques but also introduces an innovative integration of fMRI data into the image generation process, enhancing the model’s ability to interpret and visualize complex neuroimaging datasets.
@@
To leverage large-scale neural representations, we utilize the fMRI foundation model developed by Thomas et al. Thomas et al. [2022], which employs a decoder-only Transformer architecture (GPT-2) Radford et al. [2019]. As specified in Section 4.1, the input data $X$, is conformed to the preprocessed structure $X\in\mathbb{R}^{t\times 1024}$, where t is the number of TRs and 1024 denotes the parcellated brain networks. The fMRI Foundation model employs self-supervised learning (SSL) on extensive fMRI datasets, including the Human Connectome Project [Van Essen et al., 2013] and datasets available through OpenNeuro [Markiewicz et al., 2021]. It assimilates both the sequential temporal fMRI information and TR embeddings, which function similarly to position embeddings [Vaswani et al., 2017] to indicate the temporal intervals between fMRI sequences. The self-supervised learning task is to predict the subsequent timepoint sequence given an input sequence $X_{t}$; thus, the model predicts $X_{t+1}$.
In this study, we processed an fMRI run, retaining the spatial and temporal features to form the input $X$, where $X\in\mathbb{R}^{5\times 1024}$. Each fMRI segment has 5 TRs. The attention is on the embedded fMRI sequence. Based on the SSL paradigm, we took the predicted masked token, which is the output $O_{t}\in\mathbb{R}^{768}$ where $t=6$. The output $O_{6}$ was taken as our encoded fMRI latent features. Please refer to Appendix A.2 for more details.
@@
In this study, we employ multi-modal contrastive learning, capitalizing on the integration of diverse data forms to enrich the feature representation of fMRI scans in a latent space. Our approach is distinct in that it encompasses not only visual data but also textual information through the inclusion of image labels. This incorporation aims to augment the semantic understanding within our model. To facilitate this, we utilize a pre-trained Vision Transformer (ViT) Dosovitskiy et al. [2020] based CLIP model Radford et al. [2021], maintaining the integrity of its parameters by freezing them throughout the training phase.
@@
$\small\mathcal{L_{\text{CLIP}}}(z_{1},z_{2})=-\frac{1}{2|B|}\sum_{i=1}^{|B|}\left[\log\frac{\exp({z}_{\text{1}_{i}}\cdot{z}_{\text{2}_{i}}/\tau)}{\sum_{j=1}^{|B|}\exp({z}_{\text{1}_{i}}\cdot{z}_{\text{2}_{j}}/\tau)}+\log\frac{\exp({z}_{\text{2}_{i}}\cdot{z}_{\text{1}_{i}}/\tau)}{\sum_{j=1}^{|B|}\exp({z}_{\text{2}_{i}}\cdot{z}_{\text{1}_{j}}/\tau)}\right],$ (1)
@@
Furthermore, we assess the efficacy of our training process through contrastive learning accuracy metrics. Detailed implementation of these methodologies and the results are provided in A.2.
@@
To enhance our contrastive learning framework, we were inspired by prompt learning [Zhou et al., 2022, Khattak et al., 2023, Yang et al., 2024] in the AI field, which is particularly effective when text labels serve as one modality in contrastive learning. Text labels, while compact, inherently contain less information than images. Prompt learning, a technique tailored for text encoders, enriches this modality by generating context that extracts more nuanced information from the text. Considering the fMRI, image, and text share the same information. We introduced a lightweight architecture called Meta-Net designed to extract features from fMRI $\mathcal{H_{\textit{fMRI}}}$ and image $\mathcal{H_{\textit{image}}}$ latent space, $\mathcal{H}\in\mathbb{R}^{768}$, and $\text{features}~\mathcal{F}\in R^{1536}$ and output prompt learnable vectors $V\in\mathbb{R}^{8}$ for the text encoder. The vectors will be concatenated with tokenized text together input to the CLIP text encoder. Unlike CLIP text and image encoders, the gradients of the Meta-Net were not frozen during the contrastive training, where the Meta-Net was continuously trained and learnable.